OpenStudy (anonymous):

Hey peeps :D whoever answers EVERYTHING in this question, I will give them a medal every time I see a question they've answered even if it sucks. The question is: Create your own unique quadratic equation in the form y = ax^2 + bx + c that opens the same direction and shares one of the x-intercepts of the graph of y = x^2 - 10x + 16. Determine the following. Explain whether the graph has a maximum or minimum. Find the vertex and x-intercepts of the graph.

OpenStudy (anonymous):

y = ax² + bx + c y = x² - 10x + 16 0 = ( x - 2) ( x -8 ) -> a =1 which open upward so your graph will have minimum. Say your unique graph shares the x- intercepts = 2, and choose a random x = 4 -> ( x - 2) ( x - 4) = x² - 6x + 8 then your graph, obviously, has x intercepts x = 2, x = 4, and vertex ( 3, -1)